# DAGLFNet: Deep Attention-Guided Global–Local Feature Fusion for Pseudo-Image Point Cloud Segmentation

### v1

@@ 3.3 Image Feature Extraction
Considering that the input image features are projections of group-level features, which leads to sparsity and blurred boundary contours, we propose a multi-branch feature extraction architecture to enhance feature representation. Following prior work [14, 33, 16], we construct the backbone using multiple convolutional blocks. The input of the $i$-th stage is denoted as $F_{p}^{i-1}$ and $F_{g}^{i-1}$, which correspond to the feature map generated by the previous stage. The proposed BasicBlock is a multi-branch residual unit consisting of three parallel branches, designed to capture features at different receptive fields. Branch 1) employs a standard $3\times 3$ convolution to extract local features; Branch 2) uses a dilated $3\times 3$ convolution (dilation=2) [34] to enlarge the receptive field and capture richer contextual information; Branch 3) focuses on edge enhancement by first reducing the channel dimension with a $1\times 1$ convolution, followed by a $3\times 3$ convolution. Subsequently, the outputs of the three branches are concatenated along the channel dimension. A $1\times 1$ convolution is then applied to fuse these features into a unified representation. Finally, the input is added to the fused output through a residual connection, followed by an activation function, enabling direct information flow and enhancing feature representation. This step can be expressed in the following form:
@@
Here, $\text{Conv}_{k\times k}$ denotes a $k\times k$ convolution, $Conv_{3\times 3}^{d=2}$ denotes a dilated convolution with a dilation rate of 2, and $\sigma(\cdot)$ represents the normalization and activation applied after the convolution.
@@ 3.4 Feature Update Module
We take point-level and group-level features as input, and to preserve feature consistency, the local group-level features are integrated into the point-level features through a feature fusion mechanism that maintains spatial relationships and enhances discriminative power. We first project the group-level feature representation $\tilde{F}_{g}^{i-1}$ into its corresponding point-level feature space $\tilde{F}_{p}^{i-1}$ according to the projection index, using the operator $\text{InFlat}(\cdot):\mathbb{R}^{H\times W\times C}\to\mathbb{R}^{N\times C}$. The resulting point-level features are then combined with both the ${F}_{p}^{i-1}$ and the depth features of the point cloud within a tailored attention framework. Unlike conventional attention mechanisms [35, 36], our approach introduces depth information as a dynamic modulation factor to adaptively refine feature weighting. This depth-guided adjustment preserves spatial coherence during feature transformation and effectively mitigates the geometric distortions that typically arise in cross-dimensional feature fusion:
$\left\{\begin{aligned} V_{map}&=Linear(\text{InFlat}(\tilde{F}_{g}^{i-1}))\\ V_{pts}&=Linear({F}_{p}^{i-1})\\ F_{fuse\_p}&=W_{map}\odot V_{map}+W_{pts}\odot V_{pts}\\ \mathcal{F}_{p}^{i}&=Linear(F_{fuse\_p})\end{aligned}\right$. (6)
The depth values depthdepth are linearly transformed to generate the query $Q$. The $\tilde{F}_{p}^{i-1}$ are projected into key and value representations, $K_{\text{map}}$ and $V_{\text{map}}$, respectively, while the $F_{p}^{i-1}$ are similarly transformed into $K_{\text{pts}}$ and $V_{\text{pts}}$. The attention weights $W_{\text{map}}$ and $W_{\text{pts}}$ are computed by measuring the similarity between the depth query $Q$ and the corresponding keys $K_{\text{map}}$ and $K_{\text{pts}}$, followed by softmax normalization. These weights are then applied to their respective value matrices and aggregated to obtain the fused representation $F_{{fuse\_p}}$. Finally, $F_{{fuse\_p}}$ undergoes a linear transformation to produce the output feature $\mathcal{F}_{p}^{i}$.
@@ 3.5 Fusion Head Module
$\mathcal{F}_{logit}=MLP(MLP(InFlat(\mathcal{F}_{g}^{out}))+\mathcal{F}_{p}^{out})+\mathcal{F}_{p}^{0}$ (10)
@@ 4 Experiments
This section evaluates the robustness of DAGLFNet and the balance between accuracy and computational efficiency. Benchmark datasets and implementation protocols are first outlined. Quantitative and qualitative experimental results demonstrate that DAGLFNet attains an optimal trade-off between performance and efficiency, highlighting distinctive advantages. Performance and efficiency are further validated on embedded platforms, confirming practical applicability. Comprehensive ablation studies elucidate the contribution of each network component.
@@ 4.1 Datasets
We conducted comprehensive evaluations on two widely used autonomous driving LiDAR datasets. SemanticKITTI [4], collected in Karlsruhe, Germany using a Velodyne HDL-64E sensor, comprises 22 sequences with annotated point clouds, where Sequences 0–7 and 9–10 are used for training, Sequence 8 for validation, and Sequences 11–21 for online testing. Each scene contains roughly 120,000 points, annotated across 28 semantic categories, with a vertical field of view spanning $25^{\circ}$ to $3^{\circ}$. nuScenes [5], captured using a 32-beam LiDAR, includes 1,000 driving scenes with dense point clouds annotated with 32 classes; for semantic segmentation, 16 categories are evaluated. Its vertical field of view ranges from $30^{\circ}$ to $10^{\circ}$. Together, these datasets provide diverse and challenging scenarios for evaluating LiDAR-based semantic segmentation methods.
@@ 4.2 Evaluation Metrics
We use mean Intersection over Union (mIoU) and mean Accuracy (mAcc) to evaluate the semantic segmentation performance on point clouds. Suppose the dataset contains multiple semantic classes, and let $TP_{cla}$, $FP_{cla}$, and $FN_{cla}$ denote the number of true positives, false positives, and false negatives for class clacla, respectively. Then, the IoU for class clacla is defined as:
@@ 4.3 Implementation Details
Although SemanticKITTI [4] and nuScenes [5] differ in the number of LiDAR beams, both datasets share a full $360^{\circ}$ horizontal field of view. The range image resolutions are set to $64\times 1024$ and $32\times 1024$ for SemanticKITTI and nuScenes, respectively. The network depth is configured as [3, 4, 6, 3] to balance representational capacity and computational efficiency. The optimization strategy employs the AdamW [37] optimizer with an initial learning rate of 0.001, while the OneCycle [38] policy is adopted to adaptively adjust the learning rate throughout the training process. The batch size is fixed at 4 to maintain a trade-off between efficiency and memory consumption. Unless otherwise specified, all experiments were conducted on a single NVIDIA RTX 4090 GPU.
@@ 4.4 Quantitative Results
We systematically compare DAGLFNet with several state-of-the-art network architectures to comprehensively evaluate performance across multiple benchmark datasets. The evaluation encompasses both accuracy and computational efficiency, demonstrating that DAGLFNet achieves an excellent balance between the two and exhibits outstanding overall capability.
Table 1: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [4] val set. All mIoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
Table 2: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [5]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
@@
Tables 1 and 2 present a comparison of our method with existing state-of-the-art approaches on the SemanticKITTI [4] and nuScenes [5] validation sets. The results demonstrate that DAGLFNet significantly outperforms previous methods across most category-level metrics. Specifically, on the SemanticKITTI [4] validation set, our method achieves an improvement of 1.1 mIoU over WaffleIron [47] and 0.6 mIoU over FARVNet [15]; on the nuScenes [5] validation set, our method achieves an improvement of 0.7 mIoU over WaffleAndRange [53] and 0.5 mIoU over FARVNet [15], attaining the best or second-best performance across multiple categories.
Figure 3 further illustrates a comparison between DAGLFNet and the baseline method across randomly selected scenarios, for both dynamic and static classes. Notably, DAGLFNet achieves substantial improvements over the baseline in both types of classes, with an improvement of up to 23% observed in the truck category. Figure 4 shows the performance comparison between DAGLFNet and the baseline method across different distance ranges. DAGLFNet consistently outperforms the baseline, achieving improvements of 9.4% at 20–30m, 4.4% at 30–40m, and 12.5% at 40–45m. At closer ranges 10–20m, both methods perform similarly, while at 45–50m, DAGLFNet still provides a slight gain of 1.8%. Point cloud density progressively decreases with distance, creating a challenging environment for accurate discrimination. Remarkably, our method maintains strong effectiveness in these sparse, long-range regions.
Figure 3: Class-wise LiDAR segmentation results of DAGLFNet and the baseline model on the test set of SemanticKITTI.
@@
Figure 5 illustrates the relationship between mIoU and inference speed (ms/scan) on the SemanticKITTI [4] validation set for several state-of-the-art point cloud semantic segmentation methods. Our method, DAGLFNet, achieves the highest 69.1% mIoU while maintaining a moderate inference speed of 45.4 ms/scan, striking a favorable balance between accuracy and efficiency. Methods such as FRNet [14] and FARVNet [15] exhibit slightly lower accuracy with comparable inference speed, whereas faster methods like CENet [17] achieve very low latency (7.6 ms/scan) but with significantly lower accuracy (61.5% mIoU). Similarly, WaffleIron [46] shows fast inference but slightly reduced precision. Overall, Ours demonstrates a well-balanced trade-off between accuracy, inference efficiency, and model complexity, making it highly suitable for real-time applications.
@@ 4.5 Qualitative Results
Figure 6 visualizes the segmentation errors in challenging scenes on the SemanticKITTI [4] validation set. Specifically, we discuss the following four scenarios: (a) for large-scale vegetation point clouds, other methods struggle to effectively extract the most discriminative features for classification, resulting in extensive segmentation errors. In contrast, DAGLFNet can accurately segment the majority of complex regions, demonstrating superior performance; (b) furthermore, in occluded regions of similar scenes, DAGLFNet can still accurately classify the vegetation within the occluded areas, demonstrating its efficiency in local feature recognition; (c) in complex intersection scenarios, where the number of categories is large and diverse, road recognition is particularly critical. However, methods such as SphereFormer [16] and WaffleIron [46] fail to correctly classify the roads, whereas DAGLFNet can accurately identify them, ensuring precise overall scene understanding; (d) the recognition of obstacles such as vehicles at intersections is equally important. However, other methods perform poorly in identifying vehicles near turning intersections, whereas DAGLFNet demonstrates more stable and reliable recognition capabilities.
@@
To further investigate the impact of segmentation errors, we additionally visualize three types of scenarios: (a) narrow sidewalk; (b) interference cases; and (c) multiple interferences including vegetation, terrain, and occlusions, as shown in Figure 7. Specifically, the performance in three representative scenarios is as follows: (a) in narrow sidewalk scenes, WaffleIron [46], FARVNet [15], and FRNet [14] misclassify sidewalks as terrain, whereas only DAGLFNet correctly segments them; (b) in cases where sidewalks partially occlude the road, FRNet [14] and FARVNet [15] incorrectly classify the road as sidewalk. Although WaffleIron [46] correctly segments most of the road, it still exhibits confusion in certain regions, splitting the same road area into road and sidewalk, whereas DAGLFNet effectively handles this challenging situation; (c) in regions where vegetation and terrain are interwoven, other methods fail to accurately distinguish between the two, while DAGLFNet achieves accurate segmentation with minimal errors.
@@
To further evaluate the performance of DAGLFNet in distant sparse regions, we compared the segmentation of vehicles and buses , as shown in Figure 8. In distant sparse point clouds, FARVNet [15] and FRNet [14] tend to confuse sparse vehicles with adjacent sparse terrain, resulting in cars being misclassified as terrain. Under distant occlusion conditions, FARVNet [15] misclassifies buses as background buildings, while FRNet [14] only partially recognizes them and still confuses them with the background. This is mainly because conventional methods do not fully exploit global context and multi-scale features when processing sparse point clouds and local information. In contrast, DAGLFNet combines local and global features, enhances contextual information, aligns multi-scale features, and integrates a depth-guided attention mechanism, enabling distant and occluded targets to be more effectively distinguished and recognized, thereby achieving higher robustness and accuracy.
@@ 4.6 Ablation Study
Method mIoU FPS Car Bicycle Motorcycle Truck Other-vehicle Person Bicyclist Motorcyclist Road Parking Sidewalk Other-ground Building Fence Vegetation Trunk Terrain Pole Traffic-sign $64\times 256$ $65.8$ $30$ $95.1$ $50.0$ $67.4$ $82.7$ $55.7$ $65.3$ $87.6$ $0.0$ $95.0$ $64.7$ $83.3$ $24.7$ $88.1$ $60.6$ $88.8$ $61.2$ $79.1$ $55.8$ $45.7$ $64\times 512$ $68.2$ $26$ $96.9$ $54.3$ $75.7$ $71.3$ $77.1$ $75.1$ $91.1$ $0.0$ $95.6$ $63.6$ $84.3$ $14.3$ $90.3$ $66.5$ $88.1$ $66.7$ $75.1$ $61.6$ $48.8$ $64\times 768$ $68.0$ $23$ $96.7$ $51.7$ $79.7$ $92.2$ $69.5$ $77.4$ $85.2$ $0.0$ $95.7$ $47.3$ $83.5$ $10.0$ $90.9$ $66.3$ $87.1$ $61.6$ $72.6$ $66.2$ $50.3$ $64\times 1024$ $69.1$ 2121 $97.4$ $58.2$ $78.0$ $89.6$ $76.6$ $80.5$ $92.3$ $0.0$ $96.0$ $50.1$ $83.7$ $0.0$ $91.3$ $68.5$ $87.9$ $69.5$ $74.1$ $66.3$ $51.0$ $64\times 1280$ $68.6$ $17$ $96.2$ $54.7$ $77.2$ $82.2$ $68.4$ $79.6$ $90.0$ $0.0$ $94.9$ $51.6$ $84.0$ $0.0$ $91.3$ $68.8$ $87.6$ $68.3$ $73.1$ $64.6$ $51.7$
4.7 Failure Cases

### v2

@@ 3.3 Image Feature Extraction
We design a MB-FE network to process projected group-level features $\mathcal{F}_{g}^{i-1}$, addressing sparsity and blurred boundaries in pseudo-image representations. Following prior work [14, 33, 16], we construct the backbone using multiple convolutional blocks. Each BasicBlock serves as a core multi-branch unit capturing features at diverse receptive fields. Branch 1) employs a standard $3\times 3$ convolution to extract local features; Branch 2) uses a dilated $3\times 3$ convolution (dilation=2) [34] to enlarge the receptive field and capture richer contextual information; Branch 3) focuses on edge enhancement by first reducing the channel dimension with a $1\times 1$ convolution, followed by a $3\times 3$ convolution. Subsequently, the outputs of the three branches are concatenated along the channel dimension and fused via $1\times 1$ convolution. The fused output is then combined with the input through a residual connection, followed by an activation function. This process can be expressed as:
@@
Here, $\text{Conv}_{k\times k}$ denotes a $k\times k$ convolution, $\text{Conv}_{3\times 3}^{d=2}$ denotes a dilated convolution with a dilation rate of 2, and $\sigma(\cdot)$ represents the normalization and activation applied after the convolution.
@@ 3.4 Feature Update Module
The input to the feature update module consists of point-level features $\mathcal{F}{p}^{i-1}$ and group-level features $\mathcal{F}{g}^{i}$, where $\mathcal{F}_{g}^{i}$ is the output of the MB-FE network from the current stage (Section 3.3). To preserve feature consistency between these representations, we integrate the local group-level features into the point-level features through a depth-guided fusion mechanism that maintains geometric relationships and enhances discriminative power. We first project the group-level feature representation $\tilde{F}_{g}^{i-1}$ into its corresponding point-level feature space $\tilde{F}_{p}^{i-1}$ according to the projection index, using the operator $\text{InFlat}(\cdot):\mathbb{R}^{H\times W\times C}\to\mathbb{R}^{N\times C}$. To enable effective cross-feature interaction in the attention mechanism, we apply linear transformations to both the projected features $\tilde{F}_{p}^{i-1}$ and the original point-level features ${F}_{p}^{i-1}$, projecting them into a unified latent space. Unlike conventional attention mechanisms [35, 36], our approach introduces depth information as a dynamic modulation factor to adaptively refine feature weighting. The depth features depthdepth are linearly transformed to generate the query $Q$. This depth-guided adjustment preserves spatial coherence during feature transformation and effectively mitigates the geometric distortions that typically arise in cross-dimensional feature fusion:
$\left\{\begin{aligned} V_{g},K_{g}&=\text{Linear}(\text{InFlat}(\tilde{F}_{g}^{i-1}))\\ V_{p},K_{p}&=\text{Linear}({F}_{p}^{i-1})\\ Q&=\text{Linear}(depth)\end{aligned}\right$. (6)
The $\tilde{F}_{g}^{i-1}$ are projected into key and value representations, $K_{\text{g}}$ and $V_{\text{g}}$, respectively, while the $F_{p}^{i-1}$ are similarly transformed into $K_{\text{p}}$ and $V_{\text{p}}$. The attention weights $W_{\text{g}}$ and $W_{\text{p}}$ are computed by measuring the similarity between the depth query $Q$ and the corresponding keys $K_{\text{g}}$ and $K_{\text{p}}$, followed by softmax normalization. These weights are then applied to their respective value matrices and aggregated to obtain the fused representation $F_{{fuse\_p}}$. Finally, $F_{{fuse\_p}}$ undergoes a linear transformation to produce the output feature $\mathcal{F}_{p}^{i}$:
$\left\{\begin{aligned} W_{g}&=\text{softmax}(Q\odot K_{g})\\ W_{p}&=\text{softmax}(Q\odot K_{p})\\ F_{fuse\_p}&=W_{g}\odot V_{g}+W_{p}\odot V_{p}\\ \mathcal{F}_{p}^{i}&=\text{Linear}(F_{fuse\_p})\end{aligned}\right$. (7)
@@ 3.5 Fusion Head Module
$\mathcal{F}_{logit}=\text{MLP}(\text{MLP}(\text{InFlat}(\mathcal{F}_{g}^{out}))+\mathcal{F}_{p}^{out})+\mathcal{F}_{p}^{0}$ (11)
@@ 4 Experiments
This section evaluates the robustness and efficiency of DAGLFNet. We first describe the benchmark datasets and implementation protocols. Subsequently, quantitative and qualitative experimental results demonstrate the effectiveness of DAGLFNet, highlighting its favorable balance between accuracy and computational efficiency. Comprehensive ablation studies elucidate the contribution of each network component.
@@ 4.1 Datasets
We conducted comprehensive evaluations on two widely used autonomous driving LiDAR datasets. SemanticKITTI [4], the dataset was collected using Velodyne HDL-64E sensors in Karlsruhe, Germany. It comprises 22 sequences with annotated point clouds. Sequences 0–7 and 9–10 are used for training, sequence 8 for validation, and sequences 11–21 for online testing. Each scene contains approximately 120,000 points annotated across 28 semantic categories. The vertical field of view (FoV) spans $25^{\circ}$ to $3^{\circ}$. nuScenes [5], the dataset was captured using a 32-beam LiDAR. It includes 1,000 driving scenes with densely annotated point clouds across 32 classes, of which 16 categories are evaluated for semantic segmentation. The vertical field of view (FoV) ranges from $30^{\circ}$ to $10^{\circ}$. These datasets provide diverse and challenging scenarios for evaluating LiDAR-based semantic segmentation methods.
@@ 4.2 Evaluation Metrics
Segmentation performance was evaluated using mean Intersection over Union (mIoU) and mean Accuracy (mAcc). For a semantic class clacla, True Positives ($TP_{cla}$), False Positives ($FP_{cla}$), and False Negatives ($FN_{cla}$) are defined. The IoU for class cla is calculated as:
@@ 4.3 Implementation Details
Although SemanticKITTI [4] and nuScenes [5] differ in LiDAR beam count (64 vs. 32), both utilize a full $360^{\circ}$ horizontal field of view (FoV). The range image resolutions were set to $64\times 1024$ and $32\times 1024$ for SemanticKITTI [4] and nuScenes [5], respectively. The optimization strategy employs the AdamW [37] optimizer with an initial learning rate of 0.001, and the OneCycle [38] policy was used for adaptive learning rate scheduling. The batch size is fixed at 4 to maintain a trade-off between efficiency and memory consumption. All experiments were conducted on a single NVIDIA RTX 4090 GPU.
@@ 4.4 Quantitative Results
We systematically compare DAGLFNet with several state-of-the-art network architectures to comprehensively evaluate performance on the SemanticKITTI [4] and nuScenes [5] datasets. The evaluation encompasses both accuracy and computational efficiency, demonstrating that DAGLFNet achieves an excellent balance between the two and exhibits outstanding overall capability.
Table 1: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [4] val set. All mIoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline. refers to test-time augmentation.
@@
Table 2: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [5]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline. refers to test-time augmentation.
@@
The results presented in Tables 1 and 2 demonstrate the competitive performance of DAGLFNet. The results demonstrate that DAGLFNet significantly outperforms previous methods across most category-level metrics. On the SemanticKITTI [4] validation set, DAGLFNet achieves an mIoU of $69.1\%$ , surpassing WaffleIron [46] by $1.1\%$ and FARVNet [14] by $0.6\%$. On the nuScenes [5] validation set, it achieves $78.3\%$, outperforming WaffleAndRange [53] by $0.7\%$ and FARVNet [15] by 0.5%, ranking among the top performers across multiple categories.
The proposed GL-FFE, MB-FE, and FFDFA modules collectively achieve significant performance gains over the baseline, establishing DAGLFNet superiority in handling complex and sparse point clouds. Fig. 3 validates this overall advantage through consistent improvements across dynamic and static classes, with the truck category showing a 23% IoU gain that directly reflects the efficacy of GL-FFE and MB-FE in stabilizing local features and enhancing boundary discrimination for irregular structures. Fig. 4 further demonstrates module utility under extreme conditions, where the 12.5% mIoU gain at the 40–45m range specifically proves FFDFA effectiveness in leveraging depth information to prioritize valid structural features while suppressing noise in sparse regions. These results collectively confirm the critical role of all three modules in enabling robust segmentation across diverse challenging scenarios.
Figure 3: Class-wise LiDAR segmentation results of DAGLFNet and the baseline model on the val set of SemanticKITTI [4].
@@
Fig. 5 illustrates the relationship between mIoU and inference speed (ms/scan) on the SemanticKITTI [4] validation set for several state-of-the-art point cloud semantic segmentation methods. Our method, DAGLFNet, achieves the highest $69.1\%$ mIoU while maintaining a moderate inference speed of 45.4 ms/scan, striking an optimal balance between accuracy and efficiency. Methods like FRNet [14] and FARVNet [15] exhibit slightly lower accuracy with comparable inference speeds, whereas faster methods such as CENet [17] achieve very low latency (7.6 ms/scan) but with significantly lower accuracy ($61.5\%$ mIoU). Similarly, WaffleIron [46] shows fast inference but slightly reduced accuracy. Overall, DAGLFNet demonstrates a well-balanced trade-off between accuracy, inference efficiency, and model complexity.
@@ 4.5 Qualitative Results
Fig. 6 visualizes segmentation errors in challenging scenes from the SemanticKITTI [4] validation set. We examine four specific scenarios: (a) For large-scale vegetation point clouds, other methods struggle to extract the most discriminative features, leading to extensive segmentation errors. In contrast, DAGLFNet accurately segments most complex regions, outperforming others significantly. (b) Furthermore, in occluded areas of similar scenes, DAGLFNet maintains accurate classification of vegetation within the occluded regions, showcasing its effectiveness in local feature recognition. (c) In complex intersection scenarios—characterized by numerous and diverse categories—road recognition is critical. However, methods like SphereFormer [16] and WaffleIron [46] fail to classify roads correctly, whereas DAGLFNet accurately identifies them, ensuring precise scene understanding. (d) Obstacle recognition, particularly vehicles at intersections, is equally crucial. Other methods perform poorly when identifying vehicles near turning intersections, while DAGLFNet demonstrates more robust and reliable recognition capabilities.
@@
To further investigate segmentation errors, we visualize three challenging scenarios in Fig. 7: (a) narrow sidewalks, (b) occlusion cases, and (c) complex interactions involving vegetation, terrain, and occlusions. In these scenes, existing methods exhibit critical limitations: WaffleIron [46], FARVNet [15], and FRNet [14] misclassify sidewalks as terrain (a); FRNet [14] and FARVNet [15] incorrectly label occluded roads as sidewalks, while WaffleIron [46] inconsistently segments overlapping regions (b); and all struggle to distinguish intertwined vegetation from terrain (c). In contrast, DAGLFNet uniquely resolves these challenges, achieving correct segmentation of narrow sidewalks, robust handling of occlusion ambiguities, and precise separation of vegetation-terrain boundaries. These consistent failures across methods stem from a fundamental limitation: they process all points within projection subsets uniformly, regardless of distance. Consequently, noise from distant points dominates feature representations, obscuring critical geometric details in proximity and causing systematic misclassification.
@@
To further evaluate the performance of DAGLFNet in distant sparse regions, we compared the segmentation of vehicles and buses, as shown in Fig. 8. In distant sparse point clouds, FARVNet [15] and FRNet [14] tend to confuse sparse vehicles with adjacent sparse terrain, resulting in cars being misclassified as terrain. Under distant occlusion conditions, FARVNet [15] misclassifies buses as background buildings, while FRNet [14] only partially recognizes them and still confuses them with the background. This is mainly because conventional methods do not fully exploit global context and multi-scale features when processing sparse point clouds and local information. In contrast, DAGLFNet combines local and global features, enhances contextual information, aligns multi-scale features, and integrates a depth-guided attention mechanism, enabling distant and occluded targets to be more effectively distinguished and recognized, thereby achieving higher robustness and accuracy.
@@ 4.6 Ablation Study
Image Resolution mIoU FPS Car Bicycle Motorcycle Truck Other-vehicle Person Bicyclist Motorcyclist Road Parking Sidewalk Other-ground Building Fence Vegetation Trunk Terrain Pole Traffic-sign $64\times 256$ $65.8$ $30$ $95.1$ $50.0$ $67.4$ $82.7$ $55.7$ $65.3$ $87.6$ $0.0$ $95.0$ $64.7$ $83.3$ $24.7$ $88.1$ $60.6$ $88.8$ $61.2$ $79.1$ $55.8$ $45.7$ $64\times 512$ $68.2$ $26$ $96.9$ $54.3$ $75.7$ $71.3$ $77.1$ $75.1$ $91.1$ $0.0$ $95.6$ $63.6$ $84.3$ $14.3$ $90.3$ $66.5$ $88.1$ $66.7$ $75.1$ $61.6$ $48.8$ $64\times 768$ $68.0$ $23$ $96.7$ $51.7$ $79.7$ $92.2$ $69.5$ $77.4$ $85.2$ $0.0$ $95.7$ $47.3$ $83.5$ $10.0$ $90.9$ $66.3$ $87.1$ $61.6$ $72.6$ $66.2$ $50.3$ $64\times 1024$ $69.1$ 2121 $97.4$ $58.2$ $78.0$ $89.6$ $76.6$ $80.5$ $92.3$ $0.0$ $96.0$ $50.1$ $83.7$ $0.0$ $91.3$ $68.5$ $87.9$ $69.5$ $74.1$ $66.3$ $51.0$ $64\times 1280$ $68.6$ $17$ $96.2$ $54.7$ $77.2$ $82.2$ $68.4$ $79.6$ $90.0$ $0.0$ $94.9$ $51.6$ $84.0$ $0.0$ $91.3$ $68.8$ $87.6$ $68.3$ $73.1$ $64.6$ $51.7$